OpenStudy (anonymous):

Evaluate the expression without using a calculator. log8816

OpenStudy (anonymous):

\[\log _{8}8^{16}\]

OpenStudy (anonymous):

What does that mean? I'm only in sixth grade here people!

OpenStudy (anonymous):

Its logarithmics

OpenStudy (anonymous):

The answer is 16

OpenStudy (cathyangs):

with this...you drop the exponent. so its \[16\log_{8}8 \] then, try watching http://patrickjmt.com/solving-logarithmic-equations-example-1/ he has some good videos.

OpenStudy (anonymous):

Oh... Wow I'm stupid sometimes.

OpenStudy (anonymous):

The answer is 16 but how?

OpenStudy (anonymous):

log a a^x = x ... the functions are inverses

OpenStudy (cathyangs):

8 to the power of something is 8. The power is 1, and multiplying by 16, it's 16.

OpenStudy (anonymous):

recall f(f-1(x)) =x

OpenStudy (anonymous):

so also a^loga x = x